# Neural Conditional Probability for Inference

### v1

@@ 3 Operator Approach to Probability Modeling
$\bigg{|}\mathbb{P}[Y\in B\,|\,X\in A]-\mathbb{P}[Y\in B]-\frac{{\langle\mathbb{1}_{A},[\![{\mathsf{D}_{Y|X}}]\!]_{r}\mathbb{1}_{B}\rangle}}{\mathbb{P}[X\in A]}\bigg{|}\leq\sigma^{\star}_{r+1}\,\sqrt{\frac{\mathbb{P}[Y\in B]}{\mathbb{P}[X\in A]}}.$ (5)
Neural network model   Inspired by the above observations, to build the NCP model, we will parameterize the truncated SVD of the conditional expectation operator and then learn it. To that end, consider a set of parameters $(w_{i}^{\theta})_{i\in[d]}\subset\mathbb{R}$ and two parameterized embeddings $u^{\theta}\colon\mathcal{X}\to\mathbb{R}^{d}$ and $v^{\theta}\colon\mathcal{Y}\to\mathbb{R}^{d}$ as a dictionary of (nonlinear) functions
$u^{\theta}(x){:=}[u^{\theta}_{1}(x)\,\ldots\,u^{\theta}_{d}(x)]^{\top}\text{ and }v^{\theta}(y){:=}[v^{\theta}_{1}(y)\,\ldots\,v^{\theta}_{d}(y)]^{\top},\text{ with }\sigma^{\theta}{:=}[e^{-(w_{1}^{\theta})^{2}},\ldots,e^{-(w_{d}^{\theta})^{2}}]^{\top},$
parameterized by $\theta$ taking values in a set $\Theta$.
Noting that the NCP method can be applied to any neural architecture, we aim to learn the joint density function $p(x,y)$ in the form
$p_{\theta}(x,y):=1+\textstyle{\sum_{i\in[d]}}\sigma^{\theta}_{i}u^{\theta}_{i}(x)\,v^{\theta}_{i}(y)=1+{\langle\sigma^{\theta}\odot u^{\theta}_{i}(x),v^{\theta}_{i}(y)\rangle},$
where $\odot$ denotes element-wise product. To that end, we consider the loss $\mathcal{L}_{\gamma}(\theta):=\mathcal{L}(\theta)+\gamma\mathcal{R}(\theta)$ composed of two terms. The first term
$\mathcal{L}(\theta):=\mathbb{E}_{(X^{\prime},Y^{\prime})\sim\mu\times\nu}[[p_{\theta}(X^{\prime},Y^{\prime})-1]^{2}]-2\mathbb{E}_{(X,Y)\sim\rho}[p_{\theta}(X,Y)]-1$ (6)
essentially has been considered in HaoChen et al., (2022) in the specific context of augmentation graph in self-supervised deep learning, linked to kernel embeddings in Wang et al., (2022), and rediscovered and tested on DeepCCA tasks in Wells et al., (2024). Indeed, this loss can be written in terms of correlations between features. Namely, denoting the covariance and variance matrices by
$Cov[z,z^{\prime}]:=\mathbb{E}[(z-\mathbb{E}[z])(z^{\prime}-\mathbb{E}[z^{\prime}])^{\top}]\quad\text{ and }\quad Var[z]:=\mathbb{E}[(z-\mathbb{E}[z])(z-\mathbb{E}[z])^{\top}],$ (7)
@@
$\mathcal{L}(\theta):=\operatorname{\text{\rm tr}}\big{(}Var[\sqrt{\sigma^{\theta}}\odot u^{\theta}]\,Var[\sqrt{\sigma^{\theta}}\odot v^{\theta}]-2\,Cov[\sqrt{\sigma^{\theta}}\odot u^{\theta},\sqrt{\sigma^{\theta}}\odot v^{\theta}]\big{)}.$ (8)
If $p{=}p_{\theta}$ for some $\theta{\in}\Theta$, then the optimal loss is the $\chi^{2}$-divergence $\mathcal{L}(\theta){=}D_{\chi^{2}}(\rho\,|\,\mu\times\nu){=}{-}\sum_{i\geq 1}{\sigma^{\star}_{i}}^{2}$ and, as we show below, $\mathcal{L}(\theta)$ measures how well $p_{\theta}(x,y)-1$ approximates ${\sum_{i\in[r]}}\sigma^{\star}_{i}\,u^{\star}_{i}(x)\,v^{\star}_{i}(y)$. However, in order to obtain a useful probability model, it is of paramount importance to align the metric in the latent spaces with the metrics in the data-spaces $L^{2}_{\mu}(\mathcal{X})$ and $L^{2}_{\nu}(\mathcal{Y})$. For different reasons, a similar phenomenon has been observed in Kostic et al., (2024) where dynamical systems are learned via transfer operators. In our setting, this leads to the second term of the loss that measures how well features $u^{\theta}$ and $v^{\theta}$ span relevant subspaces in $L^{2}_{\mu}(\mathcal{X})$ and $L^{2}_{\nu}(\mathcal{Y})$, respectively. Namely, aiming $\mathbb{E}[u^{\star}_{i}(X)u^{\star}_{j}(X)]{=}\mathbb{E}[v^{\star}_{i}(Y)v^{\star}_{j}(Y)]{=}\mathbb{1}_{\{i=j\}}$, $i,j{\in}\{0,1,\ldots,r\}$ leads to
@@
Let ${\mathsf{E}_{Y|X}}\colon L^{2}_{\nu}(\mathcal{Y})\to L^{2}_{\mu}(\mathcal{X})$ be a compact operator and ${\mathsf{D}_{Y|X}}=\sum_{i\geq 1}\sigma^{\star}_{i}u^{\star}_{i}\otimes v^{\star}_{i}$ be the SVD of its deflated version. If $u^{\theta}_{i}\in L^{2}_{\mu}(\mathcal{X})$ and $v^{\theta}_{i}\in L^{2}_{\nu}(\mathcal{Y})$, for all $\theta\in\Theta$ and $i\in[d]$, then for every $\theta\in\Theta$, $\mathcal{L}_{\gamma}(\theta)\geq-\sum_{i\in[d]}\sigma_{i}^{\star 2}$. Moreover, if $\gamma>0$ and $\sigma^{\star}_{d}>\sigma^{\star}_{d+1}$, then the equality holds if and only if $(\sigma^{\theta}_{i},u^{\theta}_{i},v^{\theta}_{i})$ equals $(\sigma^{\star}_{i},u^{\star}_{i},v^{\star}_{i})$ $\rho$-a.e., up to unitary transform of singular spaces.
We provide the proof in App. B.2. In the following, we show how to learn these canonical features from data and construct approximations of the conditional probability measure.
@@ 4 Training the NCP inference method
In this section, we discuss how to train the model. Given a training dataset $\mathcal{D}_{n}=(x_{i},y_{i})_{i\in[n]}$ and networks $(u^{\theta},v^{\theta},\sigma^{\theta})$, we consider the empirical loss ^$\widehat{\mathcal{L}}_{\gamma}(\theta):=\widehat{\mathcal{L}}(\theta)+\gamma\widehat{\mathcal{R}}(\theta)$, where we replaced (8) and (9) by their empirical versions. In order to guarantee the unbiased estimation, as we show within the proof of Theorem 1, two terms of our loss can be written using two independent samples $(X,Y)$ and $(X^{\prime},Y^{\prime})$ from $\rho$ as
$\mathcal{L}(\theta){=}\mathbb{E}[L(u^{\theta}(X),u^{\theta}(X^{\prime}),v^{\theta}(Y),v^{\theta}(Y^{\prime}),\sigma^{\theta})]\text{ and }\mathcal{R}(\theta){=}\mathbb{E}[R(u^{\theta}(X),u^{\theta}(X^{\prime}),v^{\theta}(Y),v^{\theta}(Y^{\prime}))],$
where, the loss functionals $L$ and $R$ are defined for $u,u^{\prime},v,v^{\prime},s\in\mathbb{R}^{d}$ as
@@
$R(u,u^{\prime},v,v^{\prime})$ ${:=}(u^{\top}u^{\prime})^{2}{-}(u{-}u^{\prime})^{\top}(u{-}u^{\prime}){+}(v^{\top}v^{\prime})^{2}{-}(v{-}v^{\prime})^{\top}(v{-}v^{\prime}){+}2d,$ (11)
Therefore, at every epoch we take two independent batches $\mathcal{D}_{n}^{1}$ and $\mathcal{D}_{n}^{2}$ of equal size from $\mathcal{D}_{n}$, leading to the training algorithm given in Alg 1. See App. A.1 for the full discussion, where we also provide in Fig 2 App. A.2 an example of learning dynamics.
Algorithm 1 Condition density estimation procedure
@@
     Compute ^$\widehat{\mathcal{L}}(\theta)$ by averaging (10) over the batches
     Compute ^$\widehat{\mathcal{R}}(\theta)$ by averaging (11) over the batches
@@
Practical guidelines for training   In the following, we briefly report a few aspects to be kept in mind when using the NCP in practice, referring the reader to App. A for further details. First, the computational complexity of unbiased estimation of the loss for a dataset of size $n$ is $\mathcal{O}(nd)$, allowing one to seamlessly use NCP in contemporary DL settings. Second, the size of latent dimension $d$, as indicated by Theorem 1 relates to the problem’s ”difficulty” in the sense of smoothness of joint density w.r.t. its marginals. Lastly, after the training, an additional post-processing may be applied to ensure the orthogonality of features $u^{\theta}$ and $v^{\theta}$ and improve statistical accuracy of the learned model.
Performing inference with the trained NCP model   Now we explain how to extract important statistical objects from the trained model $(\widehat{u}^{\theta},\widehat{v}^{\theta},\sigma^{\theta})$. To that end, define the empirical operator
${\widehat{\mathsf{D}}^{\theta}_{Y|X}}{\colon}L^{2}_{\nu}(\mathcal{Y}){\to}L^{2}_{\mu}(\mathcal{X})\quad[{\widehat{\mathsf{D}}^{\theta}_{Y|X}}f](x){:=}\textstyle{\sum_{i\in[d]}}\,\sigma^{\theta}_{i}\,\widehat{u}^{\theta}_{i}(x)\,\widehat{\mathsf{E}}_{y}[\widehat{v}^{\theta}_{i}\,f],\quad\forall f{\in}L^{2}_{\nu}(\mathcal{Y}),\forall x{\in}\mathcal{X},$ (12)
@@ 5 Statistical guarantees
$\epsilon_{n}(\delta){:=}C\bigg{(}(c_{u}{\vee}c_{v})\frac{d\,\log(e\delta^{-1})}{n}{+}(\sigma_{\theta}(X){\vee}\sigma_{\theta}(Y))\sqrt{\frac{\log(e\delta^{-1})}{n}}\bigg{)},\;\overline{\epsilon}_{n}(\delta){:=}2\sqrt{2\frac{\log 2\delta^{-1}}{n}},$ (14)
@@
It follows easily from Assumption 1 that $\sigma^{2}_{\theta}(X){\leq}c_{u}^{2}d$ and $\sigma^{2}_{\theta}(Y){\leq}c_{v}^{2}d$ and consequently $\epsilon_{n}(\delta)\lesssim(c_{u}{\vee}c_{v})[\sqrt{d\log(e\delta^{-1})/n}{\vee}(d\log(e\delta^{-1})/n)]$.
@@
Discussion   The rate $\psi_{n}(\delta)$ in (16) is pivotal for the efficacy of our method. If we appropriately choose the latent space dimension $d$ to ensure accurate approximation ($\sigma^{\star}_{d+1}\ll 1$), achieve successful training ($\mathcal{E}_{\theta}\ll 1$), and secure a large enough sample size ($\varepsilon_{n}(\delta)\ll 1$), Theorem 2 provides assurance of accurate prediction of conditional probabilities. Indeed, (19) guarantees (up to a logarithmic factor)
@@
In App B.5, we derive statistical guarantees for the conditional expectation and covariance of $Y$.

### v2

@@ 3 Operator approach to probability modeling
$\bigg{|}\mathbb{P}[Y\in B\,|\,X\in A]-\mathbb{P}[Y\in B]-\frac{{\langle\mathbb{1}_{A},[\![{\mathsf{D}_{Y|X}}]\!]_{d}\mathbb{1}_{B}\rangle}}{\mathbb{P}[X\in A]}\bigg{|}\leq\sigma^{\star}_{d+1}\,\sqrt{\frac{\mathbb{P}[Y\in B]}{\mathbb{P}[X\in A]}}.$ (6)
Neural network model   Inspired by the above observations, to build the NCP model, we will parameterize the truncated SVD of the conditional expectation operator and then learn it. Specifically, we introduce two parameterized embeddings $u^{\theta}\colon\mathcal{X}\to\mathbb{R}^{d}$ and $v^{\theta}\colon\mathcal{Y}\to\mathbb{R}^{d}$, and the singular values parameterized by $w^{\theta}\in\mathbb{R}^{d}$, respectively given by
$u^{\theta}(x){:=}[u^{\theta}_{1}(x)\,\ldots\,u^{\theta}_{d}(x)]^{\top},\,\,v^{\theta}(y){:=}[v^{\theta}_{1}(y)\,\ldots\,v^{\theta}_{d}(y)]^{\top},\,\,\text{and}\,\,\sigma^{\theta}{:=}[e^{-(w_{1}^{\theta})^{2}},\ldots,e^{-(w_{d}^{\theta})^{2}}]^{\top},$
where the parameter $\theta$ takes values in a prescribed set $\Theta$.
We then aim to learn the joint density function $p(x,y)$ in the (separable) form
$p_{\theta}(x,y):=1+\textstyle{\sum_{i\in[d]}}\sigma^{\theta}_{i}u^{\theta}_{i}(x)\,v^{\theta}_{i}(y)=1+{\langle\sigma^{\theta}\odot u^{\theta}(x),v^{\theta}(y)\rangle},$
where $\odot$ denotes element-wise product. One of the prominent losses considered for the task of learning $p\in\mathcal{L}^{2}_{\mu\times\nu}$ is the least squares density ratio loss $\mathbb{E}_{\mu\times\nu}(p-p_{\theta})^{2}-\mathbb{E}_{\rho}p\!=\!\mathbb{E}_{\mu\times\nu}p_{\theta}^{2}-2\mathbb{E}_{\rho}p_{\theta}$, c.f. Tsai et al., (2020), also considered by HaoChen et al., (2022) in the specific context of augmentation graph in self-supervised deep learning, linked to kernel embeddings (Wang et al.,, 2022), and rediscovered and tested on DeepCCA tasks by Wells et al., (2024). Here, following the operator perspective, we use the characterization (4) of the optimal finite rank model to propose a new loss that: (1) excludes the known feature from the learning process, and (2) introduces a penalty term to enforce orthonormality of the basis functions. More precisely, our loss $\mathcal{L}_{\gamma}(\theta):=\mathcal{L}(\theta)+\gamma\mathcal{R}(\theta)$ is composed of two terms. The first one
$\mathcal{L}(\theta):=\mathbb{E}_{\mu\times\nu}p_{\theta}^{2}-2\mathbb{E}_{\rho}p_{\theta}+[\mathbb{E}_{\mu\times\nu}p_{\theta}]^{2}-\mathbb{E}_{\mu}[\mathbb{E}_{\nu}p_{\theta}]^{2}-\mathbb{E}_{\nu}[\mathbb{E}_{\mu}p_{\theta}]^{2}+2\mathbb{E}_{\mu\times\nu}p_{\theta}$ (7)
is equivalent to solving (4) with $A=\sum_{i=1}^{d}\sigma^{\theta}_{i}\,[u^{\theta}_{i}-\mathbb{E}_{\mu}u^{\theta}_{i}]\otimes[v^{\theta}_{i}-\mathbb{E}_{\nu}v^{\theta}_{i}]$ and can be written in terms of correlations between features. Namely, denoting the covariance and variance matrices by
${\rm Cov}[z,z^{\prime}]:=\mathbb{E}[(z-\mathbb{E}[z])(z^{\prime}-\mathbb{E}[z^{\prime}])^{\top}]\quad\text{ and }\quad{\rm Var}[z]:=\mathbb{E}[(z-\mathbb{E}[z])(z-\mathbb{E}[z])^{\top}],$ (8)
@@
$\mathcal{L}(\theta):=\operatorname{\text{\rm tr}}\big{(}{\rm Var}[\sqrt{\sigma^{\theta}}\odot u^{\theta}]\,{\rm Var}[\sqrt{\sigma^{\theta}}\odot v^{\theta}]-2\,{\rm Cov}[\sqrt{\sigma^{\theta}}\odot u^{\theta},\sqrt{\sigma^{\theta}}\odot v^{\theta}]\big{)}.$ (9)
If $p{=}p_{\theta}$ for some $\theta{\in}\Theta$, then the optimal loss is the $\chi^{2}$-divergence $\mathcal{L}(\theta){=}D_{\chi^{2}}(\rho\,|\,\mu\times\nu){=}{-}\sum_{i\geq 1}{\sigma^{\star}_{i}}^{2}$ and, as we show below, $\mathcal{L}(\theta)$ measures how well $p_{\theta}(x,y)-1$ approximates ${\sum_{i\in[d]}}\sigma^{\star}_{i}\,u^{\star}_{i}(x)\,v^{\star}_{i}(y)$. However, in order to obtain a useful probability model, it is of paramount importance to align the metric in the latent spaces with the metrics in the data-spaces $L^{2}_{\mu}(\mathcal{X})$ and $L^{2}_{\nu}(\mathcal{Y})$. For different reasons, a similar phenomenon has been observed in Kostic et al., (2024) where dynamical systems are learned via transfer operators. In our setting, this leads to the second term of the loss that measures how well features $u^{\theta}$ and $v^{\theta}$ span relevant subspaces in $L^{2}_{\mu}(\mathcal{X})$ and $L^{2}_{\nu}(\mathcal{Y})$, respectively. Namely, aiming $\mathbb{E}[u^{\star}_{i}(X)u^{\star}_{j}(X)]=\mathbb{E}[v^{\star}_{i}(Y)v^{\star}_{j}(Y)]=\mathbb{1}_{\{i=j\}}$, $i,j\in\{0,1,\ldots,d\}$ leads to
@@
Let ${\mathsf{E}_{Y|X}}\colon L^{2}_{\nu}(\mathcal{Y})\to L^{2}_{\mu}(\mathcal{X})$ be a compact operator and ${\mathsf{D}_{Y|X}}=\sum_{i=1}^{\infty}\sigma^{\star}_{i}u^{\star}_{i}\otimes v^{\star}_{i}$ be the SVD of its deflated version. If $u^{\theta}_{i}\in L^{2}_{\mu}(\mathcal{X})$ and $v^{\theta}_{i}\in L^{2}_{\nu}(\mathcal{Y})$, for all $\theta\in\Theta$ and $i\in[d]$, then for every $\theta\in\Theta$, $\mathcal{L}_{\gamma}(\theta)\geq-\sum_{i\in[d]}\sigma_{i}^{\star 2}$. Moreover, if $\gamma>0$ and $\sigma^{\star}_{d}>\sigma^{\star}_{d+1}$, then the equality holds if and only if $(\sigma^{\theta}_{i},u^{\theta}_{i},v^{\theta}_{i})$ equals $(\sigma^{\star}_{i},u^{\star}_{i},v^{\star}_{i})$ $\rho$-a.e., up to unitary transform of singular spaces.
We provide the proof in Appendix B.3. In the following section, we show how to learn these canonical features from data and construct approximations of the conditional probability measure.
Comparison to previous methods   NCP does not fall into any of the four categories defined by Gao and Hastie, (2022), as it does not aim to learn conditional density of $Y|X$ directly. Instead, NCP focuses on learning the operator mapping $L^{2}_{\nu}(\mathcal{Y})\to L^{2}_{\mu}(\mathcal{X})$, from which all relevant task-specific statistics can be derived without requiring retraining. This approach effectively integrates with deep representation learning to create a latent space adapted to $p(y|x)$. As a result, NCP efficiently captures the intrinsic dimension of the data, which is supported by our theoretical guarantees that depend solely on the latent space dimension (Theorem 2). In contrast, strategies designed for learning density often encounter significant limitations, such as the curse of dimensionality, potential substantial misrepresentation errors when the pre-specified function dictionary misaligns with the true distribution $p(y|x)$, and high computational complexity due to the need for retraining. Experiments confirm NCP’s capability to learn representations tailored to a wide range of data types—including manifolds, graphs, and high-dimensional distributions—without relying on predefined dictionaries. This flexibility allows NCP to outperform popular aforementioned methods.
@@ 4 Training the NCP inference method
In this section, we discuss how to train the model. Given a training dataset $\mathcal{D}_{n}=(x_{i},y_{i})_{i\in[n]}$ and networks $(u^{\theta},v^{\theta},\sigma^{\theta})$, we consider the empirical loss ^$\widehat{\mathcal{L}}_{\gamma}(\theta):=\widehat{\mathcal{L}}(\theta)+\gamma\widehat{\mathcal{R}}(\theta)$, where we replaced (9) and (10) by their empirical versions. In order to guarantee the unbiased estimation, as we show within the proof of Theorem 1, two terms of our loss can be written using two independent samples $(X,Y)$ and $(X^{\prime},Y^{\prime})$ from $\rho$ as $\mathcal{L}(\theta)\!=\!\mathbb{E}[L(u^{\theta}(X)\!-\!\mathbb{E}u^{\theta}(X),u^{\theta}(X^{\prime})\!-\!\mathbb{E}u^{\theta}(X^{\prime}),v^{\theta}(Y)\!-\!\mathbb{E}v^{\theta}(Y),v^{\theta}(Y^{\prime})\!-\!\mathbb{E}v^{\theta}(Y^{\prime}),\sigma^{\theta})]$ and $\mathcal{R}(\theta)\!=\!\mathbb{E}[R(u^{\theta}(X),u^{\theta}(X^{\prime}),v^{\theta}(Y),v^{\theta}(Y^{\prime}))]$, where the loss functionals $L$ and $R$ are defined for $u,u^{\prime},v,v^{\prime}\!\in\!\mathbb{R}^{d}$ and $s\!\in\![0,1]^{d}$ as
@@
$R(u,u^{\prime},v,v^{\prime})$ ${:=}(u^{\top}u^{\prime})^{2}{-}(u{-}u^{\prime})^{\top}(u{-}u^{\prime}){+}(v^{\top}v^{\prime})^{2}{-}(v{-}v^{\prime})^{\top}(v{-}v^{\prime}){+}2d.$ (12)
Therefore, at every epoch we take two independent batches $\mathcal{D}_{n}^{1}$ and $\mathcal{D}_{n}^{2}$ of equal size from $\mathcal{D}_{n}$, leading to Algorithm 1. See Appendix A.1 for the full discussion, and Appendix A.2, where we also provide in Figure 4 an example of learning dynamics.
Algorithm 1 Separable density learning procedure
@@
     Compute ^$\widehat{\mathcal{L}}(\theta)$ as an unbiased estimate using (9) or (11)
     Compute ^$\widehat{\mathcal{R}}(\theta)$ as an unbiased estimate using (10) or (12)
@@
Practical guidelines for training   In the following, we briefly report a few aspects to be kept in mind when using the NCP in practice, referring the reader to Appendix  A for further details. The computational complexity of loss estimation presents three distinct methodological approaches. The first method utilizes unbiased estimation via covariance calculations in (9) and (10), achieving a computational complexity of $\mathcal{O}(nd^{2})$ for a batch size $n$. An alternative approach employing U-statistics with (11) and (12) requires $\mathcal{O}(n^{2}d)$ operations per iteration, offering the estimation of the same precision. A third method involves batch averaging of (11) and (12), reducing computational complexity to $\mathcal{O}(nd)$, which enables seamless integration with contemporary deep learning frameworks, albeit potentially compromising training robustness through less accurate 4th-order moment estimations. Method selection remains contingent upon the specific problem’s computational and statistical constraints. Further, the size of latent dimension $d$, as indicated by Theorem 1 relates to the problem’s ”difficulty” in the sense of smoothness of joint density w.r.t. its marginals. Lastly, after the training, an additional post-processing may be applied to ensure the orthogonality of features $u^{\theta}$ and $v^{\theta}$ and improve statistical accuracy of the learned model.
Performing inference with the trained NCP model   We now explain how to extract important statistical objects from the trained model $(\widehat{u}^{\theta},\widehat{v}^{\theta},\sigma^{\theta})$. To this end, define the empirical operator
${\widehat{\mathsf{D}}^{\theta}_{Y|X}}{\colon}L^{2}_{\nu}(\mathcal{Y}){\to}L^{2}_{\mu}(\mathcal{X})\quad[{\widehat{\mathsf{D}}^{\theta}_{Y|X}}f](x){:=}\textstyle{\sum_{i\in[d]}}\,\sigma^{\theta}_{i}\,\widehat{u}^{\theta}_{i}(x)\,\widehat{\mathsf{E}}_{y}[\widehat{v}^{\theta}_{i}\,f],\quad f\in L^{2}_{\nu}(\mathcal{Y}),\,x\in\mathcal{X},$ (13)
@@ 5 Statistical guarantees
$\epsilon_{n}(\delta){:=}C\bigg{(}(c_{u}{\vee}c_{v})\frac{\sqrt{d}\,\log(e\delta^{-1})}{n}{+}(\sigma_{\theta}(X){\vee}\sigma_{\theta}(Y))\sqrt{\frac{\log(e\delta^{-1})}{n}}\bigg{)},\;\overline{\epsilon}_{n}(\delta){:=}2\sqrt{2\frac{\log 2\delta^{-1}}{n}},$ (15)
@@
It follows easily from Assumption 1 that $\sigma^{2}_{\theta}(X){\leq}c_{u}^{2}d$ and $\sigma^{2}_{\theta}(Y){\leq}c_{v}^{2}d$. Consequently, assuming that $n\geq(c_{u}\vee c_{v})d$, then $\epsilon_{n}(\delta)\lesssim(c_{u}{\vee}c_{v})\sqrt{d}[\sqrt{\log(e\delta^{-1})/n}{\vee}(\log(e\delta^{-1})/n)]$.
@@
Discussion   The rate $\psi_{n}(\delta)$ in (17) is pivotal for the efficacy of our method. If we appropriately choose the latent space dimension $d$ to ensure accurate approximation ($\sigma^{\star}_{d+1}\ll 1$), achieve successful training ($\mathcal{E}_{\theta}\approx\sigma^{\star}_{d+1}$), and secure a large enough sample size ($\varepsilon_{n}(\delta)\ll 1$), Theorem 2 provides assurance of accurate prediction of conditional probabilities. Indeed, (20) guarantees (up to a logarithmic factor)
@@
In App B.6, we derive statistical guarantees for the conditional expectation and covariance of $Y$.